OpenStudy (firejay5):

The four blades of a helicopter meet at right angles and are all the same length. The distance between the tips of two adjacent blades is 36ft. How long is each blade? Round your answer to the nearest tenth. Show and explain work in order to get medal for personage! :D

OpenStudy (anonymous):

 let the length of each blade=x ft. two adjacent blades are at right angles. applying pythagoras's theorem \[x ^{2}+x ^{2}=36^{2},2x ^{2}=36^{2},x ^{2}=\frac{ 36*36 }{ 2 }=36*18,x=18\sqrt{2}\]